Find the slope of the line passing through the points (-9,-1) and (-2,-1)
Answer by checkley75(3666)

SLOPE=(Y2-Y1)/(X2-X1)
SLOPE=(-1+1)/(-2+9)=0/7=0 OR THIS IS A LINE THROUGH Y=-1 WITH A SLOPE OF 0